OpenStudy (anonymous):

PLEASE HELP ASAP! MEDAL AND FAN! Find the value of 'm'. 18m - 108x = 6m, for x = 1

OpenStudy (anonymous):

Sub in 1 for x. Collect like terms and divide.

OpenStudy (anonymous):

If x = 1, then you don't even need to worry about it

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

ill fan all

OpenStudy (anonymous):

18m - 108 = 6m -6m -6m 12m - 108 = 0 +108 +108 12m=108 divide both sides by 12 m=9

OpenStudy (anonymous):

18m-108=6m -18m -18 -108 = -12m divide both sides by -12 m = 9